### v1

# The Open Journal of Astrophysics

Abstract
We present diffHOD-IA, a fully differentiable implementation of a halo occupation distribution (HOD) model that incorporates galaxy intrinsic alignments (IA). Motivated by the diffHOD framework of Horowitz et al. (2022), we create a new implementation that extends differentiable galaxy population modeling to include the orientation-dependent statistics crucial for weak gravitational lensing analyses. Our implementation combines this HOD formulation with the intrinsic alignment model of Van Alfen et al. (2024), enabling end-to-end automatic differentiation from HOD and IA parameters through to the galaxy field. We additionally extend this framework to differentiably model two-point correlation functions, including galaxy clustering and IA statistics. We validate diffHOD-IA against the reference halotools-IA implementation using the Bolshoi-Planck simulation, demonstrating excellent agreement across both 1-point and 2-point statistics. We verify the accuracy of gradients computed via automatic differentiation by comparing against finite difference estimates for both HOD and IA parameters. We present science use cases leveraging gradients in the simulations to uncover the IA parameters of a galaxy field representative of the tng300 simulation. We further use diffHOD-IA in a Hamiltonian Monte Carlo analysis and compare performance with halotools-IA and a neural-network based emulator of halotools-IA, named IAEmu. Unlike emulator-based approaches for statistics, diffHOD-IA provides differentiability at the catalog level, enabling its integration into field-level inference pipelines and extension to arbitrary summary statistics for next-generation weak lensing analyses.
keywords:
Weak Gravitational Lensing, Intrinsic Alignment, Differentiable Simulations
1 Introduction
The spatial distribution and intrinsic shapes of galaxies encode fundamental information about the formation and evolution of cosmic structure. Galaxy clustering, typically quantified through correlation functions, has long served as a primary probe of cosmological parameters and the galaxy-halo connection (Peebles, 1980; Davis and Peebles, 1983). More recently, the intrinsic alignments (IA) of galaxy shapes have emerged as both a significant systematic for weak gravitational lensing measurements and a cosmological signal in their own right (Troxel and Ishak, 2015; Joachimi and others, 2015; Kiessling and others, 2015; Blazek et al., 2019). As Stage IV surveys such as the Vera C. Rubin Observatory Legacy Survey of Space and Time (Ivezić et al., 2019), Euclid (Scaramella et al., 2022), and the Nancy Grace Roman Space Telescope (Akeson et al., 2019) come online, precise modeling of both galaxy clustering and IA is essential for extracting unbiased cosmological constraints from weak lensing data.
IA modeling has traditionally relied on analytic approaches rooted in perturbation theory, including the nonlinear alignment model (Hirata and Seljak, 2004; Bridle and King, 2007), the tidal alignment and tidal torquing (TATT) model (Blazek et al., 2015, 2019), and more recent effective field theory treatments (Vlah et al., 2020, 2021). While these analytic models provide valuable physical insight and computational efficiency, they often struggle to accurately capture nonlinear effects on small scales. Simulation-based approaches offer a complementary path forward. Fully nonlinear scales can be described with a halo model (Fortuna et al., 2021), which provides a framework for connecting galaxies to their host dark matter halos without explicit modeling of baryonic physics. This is in addition to semi analytic models, which combine perturbation theory approaches with halo-based methods (Maion et al., 2024). Hydrodynamic simulations provide direct predictions for galaxy alignments, although measurements exhibit significant variance across simulation suites depending on the sub-grid physics employed (Tenneti et al., 2016; Samuroff et al., 2021). For this reason, simulation frameworks that capture the galaxy-halo connection without strong dependence on subgrid physics can be valuable.
The halo occupation distribution (HOD) framework provides a powerful statistical description of the galaxy-halo connection (Peacock and Smith, 2000; Seljak, 2000; Berlind and Weinberg, 2002; Zheng and others, 2005; Zheng et al., 2007a). By specifying the probability that a halo of given mass hosts a certain number of galaxies, HOD models can populate dark matter halos from $N$-body simulations with realistic galaxy distributions. This approach has been widely successful in modeling galaxy clustering across a range of scales and redshifts (Zehavi and others, 2011; Coupon and others, 2012; Parejko and others, 2013). Extensions to the basic HOD framework have incorporated galaxy properties such as color, luminosity, and stellar mass (Zheng et al., 2007b; Zehavi and others, 2011), as well as secondary halo properties beyond mass (Hearin and others, 2016; Wechsler and Tinker, 2018).
Schneider and Bridle (2010) first introduced a halo model for IA, establishing the foundation for connecting galaxy alignments to their host dark matter halos. Building on this framework, Joachimi and others (2013b, a) developed simulation-based approaches that assign galaxy orientations by sampling from distributions encoding alignment with halo shapes, while Hoffmann et al. (2022) extended this methodology to incorporate alignment with the local tidal field. The halotools package (Hearin and others, 2017) provides a flexible framework for HOD modeling, which was extended by Van Alfen et al. (2024) to include IA modeling via the Dimroth–Watson distribution. This halotools-IA framework parameterizes the alignment strength of central and satellite galaxies separately, enabling joint modeling of galaxy clustering and orientation statistics.
A key limitation of the HOD is its reliance on stochastic sampling procedures that are not differentiable, restricting Bayesian inference with HOD simulations to methods like Markov Chain Monte Carlo (MCMC). This precludes the use of gradient-based optimization and inference methods like Hamiltonian Monte Carlo (HMC; Duane et al. 1987; Neal 2011), which have proven highly efficient in machine learning and increasingly in cosmological applications (see Dvorkin et al. (2022) for a review). As Stage IV survey analyses begin to require inference over larger parameter spaces to account for systematics, MCMC-based inference becomes increasingly computationally demanding in high-dimensional settings. For instance, the Dark Energy Survey Year 6 result included inference over 50 nuisance parameters, in addition to cosmological parameters (Collaboration et al., 2026). Consequently, there is a growing need for more computationally efficient (differentiable) methods that enable tractable inference over many free parameters.
One path forward is training NN-based emulators, which have exhibited significant forward modeling and inference speed-ups in modeling IA in hydrodynamic simulations (Jagvaral et al., 2022, 2023, 2024) and in the HOD framework (Pandya et al., 2025). Yet another path is through differentiable simulations, which enable the computation of gradients of observables with respect to model parameters via automatic differentiation. This enables gradient descent optimization, HMC sampling, and integration into differentiable analysis pipelines.
Horowitz et al. (2022) introduced diffHOD, a differentiable implementation of the HOD framework that employs continuous relaxations of discrete sampling distributions. By utilizing the Gumbel-Softmax trick (Jang et al., 2017; Maddison et al., 2017) for Bernoulli and Poisson distributions, diffHOD enables end-to-end gradient flow from HOD parameters through to galaxy catalogs. This approach was shown to accelerate parameter inference by orders of magnitude compared to traditional likelihood-free methods. Separately, Hearin et al. (2022) developed differentiable estimators for galaxy clustering statistics, enabling gradients to flow through two-point correlation function measurements.
In this work, we present diffHOD-IA, which extends the differentiable HOD framework to include galaxy intrinsic alignments. Our implementation combines the Zheng et al. (2007a) HOD formulation with the intrinsic alignment model of Van Alfen et al. (2024), enabling automatic differentiation from HOD and IA parameters through to the galaxy field and summary statistics. We develop a differentiable sampling procedure for the Dimroth–Watson distribution via inverse cumulative distribution function methods, and implement differentiable estimators for the galaxy position-position, position-orientation and orientation-orientation correlation functions. We validate diffHOD-IA against the reference halotools-IA implementation using the Bolshoi-Planck simulation (Klypin and others, 2016; Rodríguez-Puebla and others, 2016), demonstrating excellent agreement in galaxy number counts and two-point statistics. We verify gradient accuracy by comparison with finite difference estimates. We then demonstrate science applications including gradient-based optimization to recover IA parameters from mock observations and HMC inference that achieves substantial speedups over traditional MCMC approaches.
The structure of this paper is as follows. In Section 2, we review the halotools-IA formulation for HOD modeling with intrinsic alignments. Section 3 describes our differentiable implementation, including the relaxed sampling procedures and differentiable Dimroth–Watson sampling. Section 4 presents our differentiable correlation function estimators for IA statistics. In Section 5, we validate the accuracy of diffHOD-IA and its gradients. Section 6 demonstrates gradient-based optimization and HMC inference applications. We summarize and discuss future directions in Section 7.
2 The halotools-IA algorithm
We begin by reviewing the original halotools-IA formalism (Van Alfen et al., 2024, 2025b), which builds upon the halo occupation distribution (HOD) framework (Peacock and Smith, 2000; Seljak, 2000; Berlind and Weinberg, 2002), adopting the commonly used parameterization introduced by Zheng and others (2005). The halotools-IA model extends this to seven free parameters, with five of them governing the halo occupation and two of them governing the galaxy IA. These parameters are as follows:
$\{\underbrace{\log{M_{\text{min}}},\sigma_{\log{M}},\log{M_{0}},\log{M_{1}},\alpha}_{\text{HOD}},\underbrace{\mu_{\text{cen}},\mu_{\text{sat}}}_{\text{IA}}\}.$ (1)
The first two HOD parameters, $\log{M_{\text{min}}}$ and $\sigma_{\log{M}}$, are related to the occupation of central galaxies, while $\log{M_{0}}$, $\log{M_{1}}$, and $\alpha$ govern satellite galaxy occupation. More detailed descriptions of the HOD parameters can be found in Zheng et al. (2007a). The IA parameterization is a two-parameter family that statistically describes the alignment strength of central and satellite galaxies with respect to their host halos, as defined in Van Alfen et al. (2024).
2.1 Central Occupation
The central galaxy occupation is characterized by the minimum halo mass required to host a central galaxy, and the width of the transition around this threshold. These are described by the parameters $\log{M_{\text{min}}}$ and $\sigma_{\log{M}}$. The expected number of central galaxies in a halo of mass $M$ is given by:
$\langle N_{\mathrm{cen}}(M)\rangle=\frac{1}{2}\left[1+\operatorname{erf}\!\left(\frac{\log M-\log M_{\min}}{\sigma_{\log M}}\right)\right],$ (2)
where $\operatorname{erf}$ denotes the error function. Since the presence of a central galaxy is a binary outcome, central galaxies are assigned to host halos via a Bernoulli distribution:
$N_{\text{cen}}\sim\text{Bernoulli}(p=\langle N_{\mathrm{cen}}(M)\rangle).$
2.2 Satellite Occupation
Satellite galaxy occupation follows a power-law model, governed by the parameters $\log{M_{0}}$, $\log{M_{1}}$, and $\alpha$. The expected number of satellite galaxies in a halo of mass $M$ is defined as:
$\langle N_{\mathrm{sat}}(M)\rangle=\langle N_{\mathrm{cen}}(M)\rangle\left(\frac{M-M_{0}}{M_{1}}\right)^{\alpha},$ (3)
indicating that satellites are only present in halos that already host a central galaxy. Because multiple satellite galaxies can inhabit a single halo, they are sampled via a Poisson distribution:
$N_{\text{sat}}\sim\text{Poisson}(\lambda=\langle N_{\mathrm{sat}}(M)\rangle).$ (4)
We utilize a SubhaloPhaseSpace model for the IA in which satellite galaxies are always placed at the center of subhalos. These are preferentially placed in more massive subhalos first. In the event that the number of satellites exceeds the number of available subhalos, the satellite occupation resorts to a NFWPhaseSpace mode, where the remaining satellites are spatially distributed by sampling from a Navarro-Frenk-White (NFW) distribution (Navarro et al., 1996). One can also utilize the NFWPhaseSpace for all satellite occupation, in the event that subhalo information is not available.
2.3 Galaxy Intrinsic Alignments
Both galaxies and (sub)halos are modeled as triaxial homologous ellipsoids, meaning the orientations of the halos and galaxies can be described entirely from their axes. The misalignment angle of the galaxies with their host halos, $\theta_{\text{MA}}$, is governed by sampling from a Dimroth–Watson distribution (Watson, 1965). The Dimroth–Watson distribution is chosen to model galaxy alignments as it provides a maximum entropy distribution over a sphere, while accounting for the spin-2 symmetry of galaxy orientations. The probability distribution function (PDF) for the Dimroth–Watson $P(\theta,\phi)$ is defined as
$P(\theta,\phi)=\frac{B(\kappa)}{2\pi}\,e^{-\kappa\cos^{2}(\theta)}\sin(\theta)\,d\theta\,d\phi,$ (5)
for polar angle $\theta=\theta_{\text{MA}}$ and azimuthal angle $\phi$. The normalization factor is given by
$B(\kappa)=\frac{1}{2}\int_{0}^{1}e^{-\kappa t^{2}}\,dt.$ (6)
In this formulation, $\phi$ is sampled from a uniform distribution.
The fundamental parameter governing galaxy and (sub)halo alignment is $\kappa$. It is convenient to reparameterize this as
$\mu=\frac{-2\tan^{-1}(\kappa)}{\pi},$ (7)
such that $\mu=\pm 1$ corresponds to perfect (mis)alignment, and $\mu=0$ corresponds to random alignments. Both central and satellite galaxies are assigned their own alignment strengths, $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$, respectively. halotools-IA includes two separate alignment formulations for satellites: subhalo alignment, where satellites are oriented with respect to their host subhalo, and radial alignment, where satellites are oriented with respect to the host halo radial vector. The radial alignment model also admits two alignment strength possibilities: constant alignment strength, where it is the same for all galaxies, and distance-dependent alignment strength, where it has a power-law dependence on the distance to the central galaxy. More generally, the alignment strength for each galaxy can be determined based on galaxy properties. We implement all aforementioned cases in the diffHOD-IA code; in this paper, all experiments utilize the radial alignment model with constant alignment strength.
3 Differentiable halotools-IA
Figure 1: Validation of diffHOD-IA against the reference halotools-IA implementation for the tng300 fiducial HOD. Top left and center: Projected galaxy density fields across the simulation volume along the line of sight. Both implementations produce visually indistinguishable large-scale structure. Top right: Distribution of galaxy number counts $N_{\text{gal}}$ across 100 realizations using identical random seeds. Both implementations produce consistent galaxy number densities with similar scatter. Bottom left: Galaxy position-position correlation function $\xi(r)$ averaged over 100 realizations, with error bars indicating the standard deviation across realizations. The two implementations show excellent agreement across all scales. Bottom center and right: Galaxy position-orientation correlation function $\omega(r)$ and orientation-orientation correlation function $\eta(r)$. These correlations show strong agreement between the two implementations across all scales. $\eta(r)$ exhibits larger statistical noise and error bars due to the effects of galaxy shape noise. Despite the noise, the two implementations remain consistent within uncertainties.
We now proceed to outline the differentiable HOD methodology that includes galaxy IA, henceforth referred to as diffHOD-IA. In particular, diffHOD-IA allows an end-to-end differentiable mapping for:
$\text{HOD}\;+\text{IA parameters}$ $\rightarrow\text{misaligned galaxy field}$
$\rightarrow\text{summary statistics}$
under the HOD model of Zheng et al. (2007a) and IA model of Van Alfen et al. (2024). Our contributions also include differentiable modeling of two-point correlation functions (2PCFs) within this framework, following the work of Hearin et al. (2022). Differentiable computation of summary statistics is a priori nontrivial, because common cosmological summaries such as the 2PCF rely on inherently discrete operations like galaxy pair counting. We will showcase the utility of diffHOD-IA with both 2PCFs and other objectives.
For galaxy clustering, our differentiable HOD methodology closely follows that of diffHOD (Horowitz et al., 2022), which we will summarize below. diffHOD was shown to agree with (non-differentiable) HOD-generated catalogs at the level of the power spectrum to within $\sim 2.5\%$ at the scales considered ($k<3\;h/\text{Mpc}$). We extend this framework by introducing a differentiable procedure for Dimroth–Watson sampling, employing inverse cumulative distribution function (CDF) sampling. As we adopt their methodology closely, we refrain from extensively benchmarking the HOD component of diffHOD-IA, and instead focus our analyses on the IA modeling and its accuracy compared to halotools-IA.
3.1 Differentiable Sampling
Sampling from distributions is not an operation whose gradients can be tracked, as an individual sample $z$ does not encode parametric information about the distribution which it was sampled from. A common approach for backpropagating through distributions is via the reparameterization trick, as is extensively used in variational autoencoders (Kingma and Welling, 2022). In this procedure, the random variable is expressed as a deterministic function of both the distribution parameters and a source of parameter-free noise. For example, for a normally distributed variable $z\sim\mathcal{N}(\mu,\sigma^{2})$, one instead samples $\epsilon\sim\mathcal{N}(0,1)$ and rewrites $z$ as:
$z=\mu+\sigma\cdot\epsilon,$ (8)
which allows gradients to be backpropagated through $\mu$ and $\sigma$.
Discrete distributions, such as the Bernoulli or Poisson distribution, assign binary values (e.g., galaxy/no galaxy) to their random variables. As a result, it is difficult to differentiably sample from such distributions. One approach to this challenge is the Gumbel-Softmax trick (Jang et al., 2017), which defines a continuous relaxation of a discrete distribution, allowing gradients to be backpropagated via the reparameterization trick.
Specifically, for a categorical distribution with class probabilities $\{\pi_{1},\dots,\pi_{k}\}$, a sample $z$ is typically drawn as a one-hot vector using:
$z\sim\mathrm{Categorical}(\pi_{1},\dots,\pi_{k}),$ (9)
where $z_{i}=1$ indicates the selected category. However, this process is non-differentiable due to the discrete $\mathrm{argmax}$ operation implicit in categorical sampling and the inherently stochastic nature of sampling. To enable differentiability, one introduces the reparameterization trick by injecting independent Gumbel noise $g_{i}\sim\mathrm{Gumbel}(0,1)$—a choice that follows from the Gumbel-Max trick for sampling from categorical distributions (Maddison et al., 2017)—and replacing the $\mathrm{argmax}$ with a differentiable $\mathrm{softmax}$ approximation. The relaxed sample $y_{i}$ is then given by:
$y_{i}=\frac{\exp\left((\log\pi_{i}+g_{i})/\tau\right)}{\sum_{j=1}^{k}\exp\left((\log\pi_{j}+g_{j})/\tau\right)},$ (10)
where $\tau>0$ is a temperature parameter that controls the degree of approximation. In the limit $\tau\to 0$, the softmax converges to a hard (non-differentiable) categorical sample; for larger $\tau$, the distribution becomes smoother and more uniform. The trade off is in the gradients, where small $\tau$ values result in a large variance of gradients, while large $\tau$ results in a smaller variance.
3.2 Differentiable Central Occupation
Central occupation sampling is defined by a Bernoulli distribution, as described in Section 2.1. We must differentiably sample
$N_{\text{cen}}\sim\text{Bernoulli}(p=\langle N_{\mathrm{cen}}(M\;|\;M_{\text{min}},\sigma_{\log{M}})\rangle).$ (11)
To this end, we utilize the Gumbel-Softmax trick in defining the Relaxed Bernoulli distribution:
$N_{\mathrm{cen}}=\frac{1}{1+\exp\left(-\left(\log\left(\frac{p}{1-p}\right)+\epsilon\right)/\tau\right)}$ (12)
with $\epsilon\sim\mathrm{Logistic}(0,1)$. We adopt a temperature value of $\tau=0.1$, consistent with the full analysis of the occupation accuracy dependent on values of $\tau$ in Horowitz et al. (2022).
3.3 Differentiable Satellite Occupation
Satellite occupation requires sampling from a Poisson distribution as
$N_{\text{sat}}\sim\text{Poisson}(\lambda=\langle N_{\mathrm{sat}}(M\;|\;M_{0},M_{1},\alpha)\rangle).$ (13)
In Horowitz et al. (2022), it was proposed to treat each potential satellite galaxy assignment as an independently sampled Bernoulli distribution with probability $p=\lambda/N_{\text{max}}$, where $N_{\text{max}}$ is the number of trials and $\lambda$ is the Poisson rate. The resulting statistics are then Binomial distributed, which has an identical mean to a Poisson distribution. The two distributions only differ in the variance, where
$\mathrm{Var}(N_{\mathrm{sat}}^{\mathrm{Pois.}})$ $=\langle N_{\mathrm{sat}}\rangle$ (14)
$\mathrm{Var}(N_{\mathrm{sat}}^{\mathrm{Bin.}})$ $=\langle N_{\mathrm{sat}}\rangle\left(1-\frac{\langle N_{\mathrm{sat}}\rangle}{N_{\text{max}}}\right).$ (15)
In the limit $N_{\text{max}}\to\infty$, the two distributions are identical in their first two moments. In this work, we use a fiducial value of $N_{\text{max}}=48$ for all experiments, in line with the results of Horowitz et al. (2022).
We have now simplified satellite sampling into independent Bernoulli sampling identical to Equation 12 with $p=\langle N_{\mathrm{sat}}\rangle/N_{\text{max}}$. In this work, all experiments will utilize the SubhaloPhaseSpace, wherein satellite galaxies are deterministically placed at the centers of subhalos, prioritized by subhalo mass (specifically, halo_mpeak).
This prioritization is implemented differentiably by applying a softmax over the subhalos within each host halo, with logits determined by subhalo rank, i.e., the ordering of subhalos by decreasing mass within each host halo. Formally, the rank-weighted soft assignment $q_{i}$ for subhalo $i$ is computed as:
$q_{i}=\frac{\exp\left(-\mathrm{rank}_{i}/t_{\mathrm{rank}}\right)}{\sum_{j\in\mathcal{H}_{h}}\exp\left(-\mathrm{rank}_{j}/t_{\mathrm{rank}}\right)},$ (16)
where $\mathrm{rank}_{i}\in\{0,1,2,\dots\}$ is the indexed position of subhalo $i$ within host halo $h$, $\mathcal{H}_{h}$ is the set of all subhalos associated with that host, and $t_{\mathrm{rank}}=0.5$ is a temperature parameter controlling the sharpness of the prioritization. These normalized weights $q_{i}$ are then scaled by the expected number of satellites $\langle N_{\mathrm{sat}}\rangle$ to yield the relaxed per-subhalo satellite probabilities.
3.4 Differentiable NFW Sampling
In the event that $N_{\text{sat}}>|\mathcal{H}_{h}|$, we assign the locations of all remaining satellite galaxies according to samples from a NFW distribution. We generate satellite positions via an inverse CDF sampling procedure, which we will also use for IA sampling in Section 3.5. Our treatment follows the approach of Horowitz et al. (2022), but rather than using their closed-form approximation based on the Lambert $W$ function, we employ a numerically stable Newton iteration to invert the CDF. Both approaches are differentiable and similarly accurate. We begin with the CDF of the NFW profile:
$P(<r)=\frac{\ln(1+cr/r_{\mathrm{vir}})-cr/r_{\mathrm{vir}}/(1+cr/r_{\mathrm{vir}})}{\ln(1+c)-c/(1+c)},$ (17)
where $c$ is the concentration parameter of the host halo and $r_{\mathrm{vir}}$ is the virial radius of the host halo. We obtain position samples by drawing $u\sim\text{Uniform}(0,1)$ and solving for $r=P^{-1}(u)$ via Newton’s method for six iterations, which was found to be sufficient.
3.5 Differentiable Galaxy Intrinsic Alignments
The central and satellite galaxy misalignments are modeled by the Dimroth–Watson distribution, defined by its PDF in Equation 5. To differentiably sample from it, we perform inverse CDF sampling, analogous to the NFW sampling procedure described in Section 3.4.
The original (unnormalized) marginal density over the misalignment polar angle $\theta\in[0,\pi]$ is given by:
$p(\theta)\propto\exp(-\kappa\cos^{2}\theta)\,\sin\theta.$ (18)
Changing variables to $t=\cos\theta\in[-1,1]$, the density becomes:
$p(t)=\frac{1}{Z(\kappa)}\exp(-\kappa t^{2}),\qquad t\in[-1,1],$ (19)
with normalization constant
$Z(\kappa)=\int_{-1}^{1}\exp(-\kappa t^{2})\,dt.$ (20)
To draw differentiable samples, we invert the CDF
$F(t\mid\kappa)=\frac{1}{Z(\kappa)}\int_{-1}^{t}\exp(-\kappa s^{2})\,ds,$ (21)
by solving for $t$ in $t=F^{-1}(u\mid\kappa)$, where $u\sim\mathrm{Uniform}(0,1)$. This yields the following piecewise expression for the CDF:
$F(t\mid\kappa)=\begin{cases}\frac{1}{2}\left[1+\dfrac{\operatorname{erf}(\sqrt{\kappa}t)}{\operatorname{erf}(\sqrt{\kappa})}\right],&\kappa>0\\ \frac{t+1}{2},&\kappa=0\\ \frac{1}{2}\left[1+\dfrac{\operatorname{erfi}(\sqrt{-\kappa}t)}{\operatorname{erfi}(\sqrt{-\kappa})}\right],&\kappa<0\end{cases}$ (22)
As we are interested in solving $t=F^{-1}(u\;|\;\kappa)$, we must invert the expression in Equation 22 for $t$. To simplify notation, we apply the transformation $u^{\prime}=2u-1$, which maps $u\sim\mathrm{Uniform}(0,1)$ to $u^{\prime}\sim\mathrm{Uniform}(-1,1)$, and drop the prime for clarity. This yields:
$\cos(\theta)=\begin{cases}\frac{1}{\sqrt{\kappa}}\,\operatorname{erf}^{-1}\left[\operatorname{erf}(\sqrt{\kappa})\cdot u\right],&\kappa>0\\ u,&\kappa=0\\ \frac{1}{\sqrt{-\kappa}}\,\operatorname{erfi}^{-1}\left[\operatorname{erfi}(\sqrt{-\kappa})\cdot u\right],&\kappa<0\end{cases}$ (23)
from which we can compute $\theta$. To construct the full 3D orientation vector $\bm{n}$, we draw $u\sim\mathrm{Uniform}(-1,1)$ to obtain $t=\cos\theta$ using Equation 23, and independently sample $\phi\sim\mathrm{Uniform}(0,2\pi)$. These are converted to Cartesian coordinates as:
$\bm{n}=\begin{bmatrix}\sin\theta\cos\phi\\ \sin\theta\sin\phi\\ \cos\theta\end{bmatrix}$. (24)
In practice, the inverse CDF is solved via Newton’s method with six iterations. The resulting galaxy orientation vectors are fully differentiable with respect to the IA parameters.
4 Differentiable Correlation Functions
Typical cosmological analyses employ low dimensional summary statistics computed over the full galaxy field. We proceed by outlining a prescription to differentiably calculate 2PCFs, enabling an end-to-end differentiable pipeline from halo occupation and IA parameters to summary statistics. We find that computing the IA statistics is directly differentiable with respect to the IA parameters using discrete galaxy catalogs; computing $\xi(r)$ requires generating catalogs where the galaxies have a weight proportional to their occupation probability, as done in Hearin et al. (2022). We note that the IA statistics $\omega(r)$ and $\eta(r)$ also depend on the HOD parameters through the galaxy positions and number counts; however, for the experiments in this work, we operate in a fixed HOD setting to validate the differentiable IA implementation. We develop and benchmark accuracy for both the weighted $\xi(r)$ and IA correlation estimators in Appendix B. All correlation measurements use 20 logarithmically-spaced bins from $r=0.1\;h^{-1}\text{Mpc}$ to $r=16\;h^{-1}\text{Mpc}$.
4.1 Intrinsic Alignment Statistics
In this context, there are two IA statistics of interest: $\omega(r)$ and $\eta(r)$. The galaxy position-orientation correlation function $\omega(r)$ is defined as:
$\omega(r)=\langle(\hat{e}(\mathbf{x})\cdot\hat{r})^{2}\rangle-\frac{1}{3},$ (25)
where $\hat{e}$ is the galaxy orientation unit vector that specifies the intrinsic orientation of each galaxy’s major axis, $\mathbf{x}$ is the position vector of a galaxy, and $\hat{r}$ is the unit vector in the direction of galaxy separation. The galaxy orientation-orientation correlation function $\eta(r)$ is defined as:
$\eta(r)=\langle(\hat{e}(\mathbf{x})\cdot\hat{e}(\mathbf{x}+\mathbf{r}))^{2}\rangle-\frac{1}{3},$ (26)
which measures the correlation between the orientations of galaxy pairs as a function of their separation.
We pre-compute all neighbor pairs $(i,j)$ with separations $|\mathbf{r}_{ij}|<r_{\rm max}$ using a KD-tree with periodic boundary conditions. This is implemented via scipy. Precomputation strategies to avoid redundant pair-finding operations have been employed previously to accelerate HOD parameter exploration, including tabulating pair counts as a function of halo mass (Neistein et al., 2011; Zheng and Guo, 2016). Our approach differs in that we store the explicit pair indices for a fixed galaxy catalog, enabling gradient flow through the orientation-dependent alignment statistics while treating the neighbor structure as constant. For each pair, we compute the separation vector with periodic wrapping. For $\omega(r)$, we compute the position-orientation alignment:
$a_{ij}^{\omega}=\left(\hat{e}_{i}\cdot\hat{r}_{ij}\right)^{2}.$ (27)
For $\eta(r)$, we compute the orientation-orientation alignment:
$a_{ij}^{\eta}=\left(\hat{e}_{i}\cdot\hat{e}_{j}\right)^{2},$ (28)
which measures the alignment between the orientation vectors of galaxies $i$ and $j$.
Pairs are assigned to radial bins, and the correlation functions are estimated as:
$\omega(r_{k})=\frac{\sum_{(i,j)\in\mathcal{B}_{k}}a_{ij}^{\omega}}{|\mathcal{B}_{k}|}-\frac{1}{3}$ (29)
and
$\eta(r_{k})=\frac{\sum_{(i,j)\in\mathcal{B}_{k}}a_{ij}^{\eta}}{|\mathcal{B}_{k}|}-\frac{1}{3}.$ (30)
Here $\mathcal{B}_{k}$ denotes the set of all galaxy pairs whose separations fall within radial bin $k$, and $|\mathcal{B}_{k}|$ is the number of such pairs. Crucially, since galaxy positions are fixed, the pair counts $|\mathcal{B}_{k}|$ are constants with respect to the IA parameters. Gradients flow exclusively through the orientation vectors $\hat{e}_{i}$, which depend on $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$ via the differentiable Dimroth–Watson sampling described in Section 3.5.
5 Accuracy & Gradients
Table 1: Bolshoi-Planck simulation parameters.
Particle Mass $\Omega_{m,0}$ $\sigma_{8}$ $n_{s}$ $h$ $L_{\rm box}$ $z$
($h^{-1}M_{\odot}$) ($h^{-1}$Mpc)
$\sim 10^{8}$ 0.30711 0.82 0.96 0.70 250 0
To evaluate diffHOD-IA, we assess the accuracy of its IA implementation by comparing it against halotools-IA. We also examine the model’s differentiability with respect to both HOD and IA parameters, and compare the resulting gradients to those obtained via finite difference methods. A detailed performance benchmark of the HOD component is not provided, as its implementation closely follows that of Horowitz et al. (2022). However, we do benchmark the computational cost of diffHOD-IA by comparing $N_{\rm max}$, used in the differentiable satellite occupation, to $N_{\rm Newton}$, used for inverse CDF sampling from the Dimroth–Watson distribution. These results are presented in Appendix A.
We will construct a mock observable galaxy catalog and use this fiducial catalog as a reference in the following sections. This catalog is constructed using the parameters:
$\log M_{\min}=12.54,\;\sigma_{\log M}=0.26,\;\log M_{0}=12.68$
$\log M_{1}=13.48,\;\alpha=1.0,\;\mu_{\text{cen}}=0.79,\;\mu_{\text{sat}}=0.30.$
The HOD parameters were determined by fitting the halo occupation to match a galaxy catalog from the tng300 (Nelson et al., 2021) simulation for a stellar mass $M_{*}$ cutoff of $\log M_{*}\geq 10.5$. This HOD is run on the Bolshoi-Planck dark matter catalog at $z=0$, whose simulation parameters can be found in Table 1. We adopt the HOD configuration used in Van Alfen et al. (2024); Pandya et al. (2025); details of how the catalog was constructed to match tng300 are given in Van Alfen et al. (2024), and the IA parameters correspond to the tng300 best-fit values reported in Pandya et al. (2025).
5.1 Comparison to halotools-IA
We compare the performance between diffHOD-IA and halotools-IA using visualizations of the galaxy field density, comparison of 1-pt statistics (i.e., galaxy number counts, $N_{\text{gal}}$), and relevant 2-pt statistics. This is shown in Figure 1. This comparison provides a comprehensive metric for the performance of diffHOD-IA.
We find excellent agreement between diffHOD-IA and halotools-IA for the fiducial galaxy catalog used. The galaxy field densities produced by the two simulations are visually consistent, as shown in the top left and center panels of Figure 1. As galaxy placements are deterministic due to the use of SubhaloPhaseSpace, this is representative of agreement in the per-halo $\langle N_{\rm cen}\rangle$ and $\langle N_{\rm sat}\rangle$ between the two implementations. In both cases, approximately $0.17\%$ of the remaining satellites were occupied according to NFWPhaseSpace.
To go beyond a visual comparison, this is further confirmed upon examining the histogram of $N_{\text{gal}}$ across 100 realizations in the top right panel of Figure 1. The diffHOD-IA galaxy number counts are $29772\pm 109$, and the halotools-IA number counts are $29729\pm 110$, illustrating excellent agreement. Minor differences in the mean values may stem from the fact that the seed mapping between diffHOD-IA and halotools-IA is not one-to-one, which can lead to small deviations over a finite number of realizations.
In the bottom panels of Figure 1, we compare $\xi(r)$, $\omega(r)$, and $\eta(r)$ between the two implementations. This is using the (non-differentiable) halotools-IA implementation for the correlation estimators, to highlight any potential differences at the 2PCF level coming solely from the galaxy occupation and IA. We find generally excellent agreement between diffHOD-IA and halotools-IA across all three statistics. This even includes the IA statistic $\eta(r)$, which is much noisier than $\omega(r)$ due to galaxy shape noise. We additionally see comparable error bars between the implementations across 100 realizations. We benchmark the accuracy of the differentiable correlation estimators as outlined in Section 4 in Appendix B.
5.2 HOD Gradients in diffHOD-IA
Figure 2: Gradients of the halo occupation distribution functions with respect to HOD parameters as a function of halo mass. Left panel: Gradients of the mean central galaxy occupation $\langle N_{\rm cen}\rangle$ with respect to $\log M_{\rm min}$ (blue) and $\sigma_{\log M}$ (orange). The gradients are largest in the transition region around $\log_{10}M/M_{\odot}\approx 12$ where the occupation probability transitions from 0 to 1, and vanish at high masses where $\langle N_{\rm cen}\rangle$ saturates to unity. Right panel: Gradients of the mean satellite galaxy occupation $\langle N_{\rm sat}\rangle$ with respect to all five HOD parameters: $\log M_{\rm min}$, $\sigma_{\log M}$, $\log M_{0}$, $\log M_{1}$, and $\alpha$. In both panels, solid lines show gradients computed via automatic differentiation and dotted points show finite difference estimates, demonstrating excellent agreement. The IA parameters $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$ do not affect galaxy number counts and have zero gradient everywhere. These gradients enable efficient gradient-based inference of HOD parameters from galaxy clustering observations.
We first study gradients of the 1-pt statistics of the galaxy catalog, $\langle N_{\text{cen}}\rangle$ and $\langle N_{\text{sat}}\rangle$, for various halo masses evaluated at the fiducial HOD values. We compare the autodifferentiation (autodiff) values from diffHOD-IA with finite-difference methods. The IA parameters, $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$, are excluded in this analysis as they do not affect the galaxy number counts. The results of this analysis are shown in Figure 2.
We see in the left panel of Figure 2 that the only HOD parameters with non-zero gradients for $\langle N_{\text{cen}}\rangle$ are $\log M_{\text{min}}$ and $\sigma_{\log M}$, which is in agreement with the analytic expression for $\langle N_{\text{cen}}\rangle$ given in Equation 2. The finite difference gradients are additionally plotted as scatter, exhibiting excellent agreement with the autodiff gradients. The gradients vanish for $\log_{10}M/M_{\odot}\gtrsim 13$, at which point all available halos in the catalog are populated with a central galaxy. Similarly, at $\log_{10}M/M_{\odot}\lesssim 11$, the halos are not massive enough to frequently host central galaxies, so the gradients become small. The gradients for both $\sigma_{\log M}$ and $\log M_{\text{min}}$ are approximately unity (up to a sign) at $\log_{10}M/M_{\odot}\approx 12$, where the occupation probability for the fiducial HOD transitions from 0 to 1.
In the right panel of Figure 2, we see that all five HOD parameters have nonzero gradients for $\langle N_{\text{sat}}\rangle$, as expected from the analytic expression given in Equation 3. The magnitudes of the gradients vary across several orders of magnitude as halo mass increases, which is an important diagnostic for specifying learning rates when using the gradients in a gradient-based optimization pipeline. We again see excellent agreement between the diffHOD-IA autodiff gradients and the finite differences. These results are in excellent agreement with a similar analysis shown in Horowitz et al. (2022).
5.3 Intrinsic Alignment Gradients in diffHOD-IA
Figure 3: Validation of differentiable sampling from the Dimroth–Watson distribution for galaxy-halo misalignment angles. Left panel: Probability distribution $P(\cos\theta)$ of misalignment angles for varying alignment strength $\mu$. The top axes show the corresponding misalignment angle $\theta$ in degrees. Solid lines show histograms from samples drawn using our differentiable inverse-CDF sampler; dashed lines show the analytic Dimroth–Watson PDF. The close agreement validates the differentiable sampling implementation. Positive $\mu$ (magenta) produces alignment with probability concentrated at $\cos\theta=\pm 1$, while negative $\mu$ (cyan) produces anti-alignment peaked at $\cos\theta=0$. Right panel: Gradient of the probability distribution with respect to the alignment parameter, $\partial P/\partial\mu$. Dashed lines show analytic gradients derived from the PDF formula; scatter points show finite-difference gradients of the analytic Dimroth–Watson PDF; solid lines show gradients computed via automatic differentiation through the sampling procedure, where a Gaussian kernel density estimate is used to obtain a smooth density from discrete samples.
We next examine the differentiability of the IA model in diffHOD-IA by validating the accuracy and gradients of the Dimroth–Watson distribution sampling procedure. The Dimroth–Watson distribution is parameterized by the alignment strength $\mu\in[-1,1]$, which controls the shape of the Dimroth–Watson distribution from which galaxy misalignment angles are sampled. We compare the differentiable Dimroth–Watson samples with typical samples for a range of misalignment angles in Figure 3. The diffHOD-IA data was generated using $500000$ samples and 6 Newton iterations for the inverse CDF sampling procedure. For each sample, we compute the misalignment angle $\theta$ between the sampled galaxy orientation and a reference alignment direction.
The left panel of Figure 3 shows the empirical probability distributions $P(\cos\theta)$ constructed from the samples (solid lines) alongside the analytic Dimroth–Watson PDF (dashed lines). The histograms are computed with 100 bins spanning $\cos\theta\in[-1,1]$ and normalized to unit area. The close agreement across all values of $\mu$ validates our implementation of the inverse-CDF sampling procedure for sampling from the Dimroth–Watson. As expected, positive $\mu$ produces alignment with probability mass concentrated near $\cos\theta=\pm 1$ (corresponding to $\theta\approx 0^{\circ}$ or $180^{\circ}$), while negative $\mu$ produces anti-alignment with probability peaked at $\cos\theta=0$ ($\theta\approx 90^{\circ}$).
To validate the gradients, we compute $\partial P(\cos\theta)/\partial\mu$ in three ways. First, we use autodiff through the diffHOD-IA sampling procedure. We draw samples for each $\mu$ value, construct a differentiable Gaussian kernel density estimate (KDE) to obtain a smooth density function from the discrete samples, and compute gradients via autodiff. The KDE is necessary to convert the discrete histogram into a continuous, differentiable function. We additionally compute analytic and finite-difference gradients of the exact Dimroth–Watson PDF with respect to $\mu$, which can be derived from Equation 5 for the analytic case. The right panel of Figure 3 shows generally good agreement between the autodiff gradients (solid lines), the analytic gradients (dashed lines), and finite-difference gradients (scatter). We note that in the case of $\mu\approx 0$, the gradient similarly tends to zero as the Dimroth–Watson distribution becomes uniform. The gradient can thus be sensitive to Monte Carlo noise in this regime.
6 Applications of differentiable capabilities
We demonstrate the utility of diffHOD-IA with gradient-based optimization experiments. These experiments illustrate how gradients in the simulation can be used to obtain parameter estimates and posteriors over parameters of interest from mock observational data. All experimental results presented use a value of $N_{\text{max}}=48$, $\tau=0.1$, and SubhaloPhaseSpace with a NFWPhaseSpace fallback. A constant radial alignment strength model is used for the subhalo IA. As our construction follows Horowitz et al. (2022) closely, and having validated both the accuracy and differentiability of diffHOD-IA against halotools-IA, our emphasis will be on the IA parameters $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$; several experiments illustrating the differentiability of the HOD are included in Horowitz et al. (2022). We construct moment-matching and differentiable correlation function optimization objectives. In addition, we demonstrate accelerated inference with diffHOD-IA using HMC, compared to halotools-IA with MCMC.
6.1 Moment-matching Objective
Setup. We consider the task of inferring the IA parameters $\bm{\mu}=(\mu_{\text{cen}},\mu_{\text{sat}})$ from a target galaxy catalog observable, restricting for simplicity to a single HOD configuration given by the fiducial tng300 model used previously. Importantly, fixing the HOD does not fix the realized galaxy catalog: different random seeds produce catalogs with varying numbers of galaxies, requiring an optimization procedure that is robust to stochastic catalog realizations. To this end, we consider a moment matching optimization procedure, matching the misalignment angle distributions between the generated and target galaxy catalogs via their first two moments. This does not rely on a one-to-one correspondence between galaxies, and additionally requires no knowledge of the underlying PDF governing the galaxy misalignments.
Optimization. For a given parameter vector $\bm{\mu}$ and simulation seed $s$, we generate the mock galaxy catalog and compute the per-galaxy alignment statistic $t^{2}=(\mathbf{n}\cdot\mathbf{u})^{2}$, where $\mathbf{n}$ is the galaxy orientation and $\mathbf{u}$ is the reference axis (host halo major axis for centrals, radial direction for satellites). We use $t^{2}$ rather than $t$ because the Dimroth–Watson distribution is symmetric about $t=0$, making $\langle t\rangle=0$ for all $\mu$. For each population (centrals and satellites separately), we compute the mean $\langle t^{2}\rangle$ and variance $\mathrm{Var}(t^{2})$ across all galaxies, which serve as our summary statistics.
Our loss function matches these first and second moments between simulated and observed catalogs across seeds $s$, treating central and satellite galaxy populations separately
$\mathcal{L}(\bm{\mu};s)=\alpha_{\text{cen}}\mathcal{L}_{\text{cen}}+\mathcal{L}_{\text{sat}},$ (31)
where
$\mathcal{L}_{\text{cen}}$ $=\left(\langle t^{2}\rangle_{\mathrm{sim}}-\langle t^{2}\rangle_{\mathrm{obs}}\right)^{2}$
$\quad+w_{\text{cen},\sigma^{2}}\left(\mathrm{Var}[t^{2}]_{\mathrm{sim}}-\mathrm{Var}[t^{2}]_{\mathrm{obs}}\right)^{2},$ (32)
$\mathcal{L}_{\text{sat}}$ $=\left(\langle t^{2}\rangle_{\mathrm{sim}}-\langle t^{2}\rangle_{\mathrm{obs}}\right)^{2}$
$\quad+w_{\text{sat},\sigma^{2}}\left(\mathrm{Var}[t^{2}]_{\mathrm{sim}}-\mathrm{Var}[t^{2}]_{\mathrm{obs}}\right)^{2},$ (33)
where it is implied that the $t^{2}$ statistics are computed over central galaxies only in $\mathcal{L}_{\text{cen}}$ and satellites only in $\mathcal{L}_{\text{sat}}$. We use weights $w_{\text{cen},\sigma^{2}}=w_{\text{sat},\sigma^{2}}=0.5$, and $\alpha_{\text{cen}}=2.0$ to up-weight the smaller central galaxy population. To mitigate the Monte Carlo noise in Dimroth–Watson sampling, we average the loss over $N_{s}=3$ random seeds:
$\bar{\mathcal{L}}(\bm{\mu})=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\mathcal{L}(\bm{\mu};s_{i}).$ (34)
This construction matches the variance $\mathrm{Var}[t^{2}]$ within each catalog to capture galaxy shape noise, while averaging the loss over multiple simulated realizations stabilizes the gradient signal. The seed averaging is performed over a fixed set of seeds; importantly, the target catalog uses a different seed than the optimization seeds. We compare experiments using one and three seeds in Figure 4. More detail on optimization hyperparameters is given in Appendix C.
Figure 4: Gradient-based recovery of IA parameters from 50 random initializations using moment-matching optimization. The target parameters ($\mu_{\rm cen}=0.7905$, $\mu_{\rm sat}=0.307$, pink star) represent the best-fit $\mu$ values to the tng300 HOD configuration, with the empirical uncertainty shown as pink contours. Optimization trajectories are shown as gray lines connecting initial positions to converged solutions. Brown circles denote optimizations using a single HOD realization seed per gradient step, while green squares show results when averaging over three seeds. The inset panel (lower left) shows a zoomed view of the convergence region, revealing tight clustering of final parameter estimates around the true values. Both single-seed and three-seed strategies successfully recover the target parameters across diverse initializations.
We note that for the target catalog generated at fixed input parameters, the observed alignment angles represent a single Monte Carlo draw from the underlying Dimroth–Watson distribution, inducing an effective $\bm{\mu}$ that may differ slightly from the nominal input values. To determine the effective $\bm{\mu}$ values for the target catalog, we independently match the observed $\langle t^{2}\rangle$ to the theoretical Dimroth–Watson expectation for the central and satellite populations. This yields best-fit values of $\mu_{\text{cen}}=0.7905$ and $\mu_{\text{sat}}=0.307$, which differ slightly from the input values $\mu_{\text{cen}}=0.79$ and $\mu_{\text{sat}}=0.30$.
We estimate the uncertainties in $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$ by quantifying the variance in effective $\mu$ values across independent catalog realizations. To this end, we generate 50 independent HOD realizations at the fiducial IA parameters using different random seeds. For each realization, we compute $\langle t^{2}\rangle$ separately for centrals and satellites, as done in the optimization, and determine the effective $\mu$ by matching to the theoretical Dimroth–Watson expectation via numerical inversion. The scatter in these effective $\mu$ values across realizations quantifies the uncertainty in the effective $\mu$ for a single catalog realization, capturing the combined variance from HOD sampling and orientation sampling. Importantly, this also defines the expected region within which our optimization should converge. This yields uncertainties $\sigma(\mu_{\text{cen}})=0.001$ and $\sigma(\mu_{\text{sat}})=0.012$, with a correlation coefficient of $\rho=0.04$. These uncertainties are visualized in pink in Figure 4.
Results. Figure 4 shows the results of gradient-based optimization of IA parameters using the moment matching objective. We perform 50 independent optimization runs from random initializations uniformly sampled from $\mu\in[-1.0,1.0]$, each run for 100 optimization steps. The trajectories in Figure 4 show the parameter space flow from initializations to the fiducial tng300 IA values, with final points filtered by best loss. We see that the moment-matching loss over the full galaxy field successfully optimizes the IA parameters toward the ground truth values for the HOD across all initializations and both seed-averaging strategies. In the inset panel, we observe that the 3-seed optimization exhibits slightly tighter convergence to the true values. Quantitatively, the 1-seed optimization results in final values $\mu_{\text{cen}}=0.791\pm 0.002$ and $\mu_{\text{sat}}=0.320\pm 0.007$, while the 3-seed optimization achieves $\mu_{\text{cen}}=0.790\pm 0.002$ and $\mu_{\text{sat}}=0.309\pm 0.006$.
The scatter in converged $\mu_{\text{cen}}$ values is slightly larger than the computed target uncertainty, likely reflecting optimization noise and finite optimization time. Conversely, the scatter in converged $\mu_{\text{sat}}$ values is comparable to the target uncertainty, and the slight bias toward higher values may arise from an asymmetry of the loss landscape near the true value. Nevertheless, all optimization runs converge to within a few percent of the true values, demonstrating robust parameter recovery. Although the IA parameters in this simplified setting could be recovered via a direct curve fit, this experiment illustrates how differentiability enables efficient parameter recovery using only one-point alignment statistics. This is especially relevant when the analytic form of the PDF is not known; in this case, we have captured properties of the true distribution as characterized by its first two moments.
6.2 Correlation Function Objective
Figure 5: Gradient-based recovery of IA parameters from 50 random initializations using correlation function matching optimization. The target parameters ($\mu_{\rm cen}=0.79$, $\mu_{\rm sat}=0.30$, pink star) represent the fiducial tng300 HOD configuration, with $1\sigma$ and $2\sigma$ uncertainty regions estimated from the scatter of maximum-likelihood fits across 50 independent HOD realizations shown in pink. Optimization trajectories are shown as gray lines connecting initial positions to converged solutions. Gray circles denote converged values. The inset panel (lower left) shows a zoomed view of the convergence region, revealing tight clustering of final parameter estimates around the true values.
Setup. We now turn to optimization with a more traditional summary statistic with 2PCFs. In this section, we will optimize the IA parameters such that the generated catalog $\omega(r)$ correlation matches that of the fiducial tng300 correlation function, denoted $\hat{\omega}(r)$. $\hat{\omega}(r)$ is computed by averaging over 20 independent orientation samples at the fiducial IA parameters. The per-bin variance of $\hat{\omega}(r)$ is estimated from the scatter across these realizations, which reduces galaxy shape noise in the optimization target. Unlike a single realization of galaxy orientations, which can result in effective $\bm{\mu}$ values that differ from the input, this averaging ensures that the optimal $\bm{\mu}$ parameters for $\hat{\omega}(r)$ converge to the true input values. We choose to optimize on this statistic over $\eta(r)$, as $\omega(r)$ is less contaminated with galaxy shape noise due to it being a cross correlation with galaxy positions. We cannot use $\xi(r)$ as it has no dependence on $\bm{\mu}$; a full analysis that includes the HOD parameters would jointly use $\xi(r)$, $\omega(r)$, and $\eta(r)$.
Optimization. We construct a weighted mean-squared-error loss between the predicted and target correlation functions:
$\mathcal{L}(\bm{\mu})=\sum_{k}W_{k}\left[\omega(r_{k})-\hat{\omega}(r_{k})\right]^{2},$ (35)
where $\hat{\omega}(r_{k})$ is the averaged target correlation computed from the fiducial tng300 catalog and $W_{k}$ are inverse-variance weights, determined by estimating the variance of $\hat{\omega}(r_{k})$ across multiple orientation realizations at fixed positions. We generate $20$ independent orientation samples at reference IA parameter values and compute:
$W_{k}=\frac{1}{\sigma^{2}_{\omega}(r_{k})+\epsilon},$ (36)
where $\sigma^{2}_{\omega}(r_{k})$ is the empirical variance in bin $k$ and $\epsilon=10^{-6}$ provides numerical stability to prevent division by zero in bins with very low variance. The weights are normalized such that $\sum_{k}W_{k}=1$. This inverse-variance weighting approximates the covariance due to galaxy shape noise, which is in general dominant over the sample variance at the scales being considered (Van Alfen et al., 2024). While the jackknife covariances that account for both sample variance and galaxy shape noise from the TNG300 data itself are available following the procedure of Van Alfen et al. (2024), we instead use a diagonal variance estimate obtained by averaging over many diffHOD-IA orientation realizations (for fixed galaxy positions), which serves as a proxy for the shape noise and is considerably more computationally efficient. More detail on optimization hyperparameters is given in Appendix C.
Results. Figure 5 shows the results of gradient-based optimization of IA parameters using the differentiable $\omega(r)$ objective. We perform 50 independent optimization runs from random initializations uniformly sampled from $\mu\in[-1.0,1.0]$, each run for 2000 optimization steps. The optimization successfully recovers the target parameters across all initializations, with mean recovered values $\mu_{\text{cen}}=0.791\pm 0.003$ and $\mu_{\text{sat}}=0.303\pm 0.009$, compared to target values of $\mu_{\text{cen}}=0.79$ and $\mu_{\text{sat}}=0.30$. All converged solutions lie well within the expected statistical uncertainty, shown in pink in Figure 5.
To estimate the uncertainty, we generate 50 independent HOD catalog realizations at the true IA parameters with different random seeds. For each realization, we compute $\omega(r)$ averaged over 20 orientation samples and find the maximum likelihood IA parameters via grid search, minimizing
[…]
where $\mathbf{C}$ is the covariance matrix of $\omega(r)$ estimated from 50 independent orientation samples at the fiducial parameters. We additionally correct for the finite number of realizations by incorporating the Hartlap factor (Hartlap et al., 2006) when inverting the covariance matrix. The scatter in recovered parameters across these 50 catalogs provides an empirical estimate of the covariance, yielding uncertainties $\sigma(\mu_{\text{cen}})=0.008$ and $\sigma(\mu_{\text{sat}})=0.026$, with correlation coefficient $\rho=-0.68$. This anti-correlation aligns with theoretical expectations and was also seen experimentally in Berman et al. (2025). The tighter clustering of converged values reflects the fact that all optimization runs fit the same catalog to the same target, whereas the uncertainty contours capture the full variance including HOD stochasticity.
6.3 Hamiltonian Monte Carlo
Figure 6: $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$ posteriors for the fiducial tng300 catalog derived using diffHOD-IA with HMC (blue), IAEmu with HMC (orange), and halotools-IA with MCMC (green). diffHOD-IA is in excellent agreement with halotools-IA, while exhibiting substantially faster inference.
Setup. To complement our previous correlation function matching gradient descent experiment, we now turn to HMC to illustrate Monte Carlo inference capabilities with diffHOD-IA. We use the same $\hat{\omega}(r)$ mock observation from the previous section, this time using the jackknife covariance estimate from tng300. For this analysis, we use a fixed HOD and only consider $\mu_{\text{sat}}$ and $\mu_{\text{cen}}$ as free parameters.
We use the uninformative priors:
$\mu_{\text{cen}},\mu_{\text{sat}}\sim\text{Uniform}(-1,1),$ (38)
and employ the $\tt{numpyro}$ (Phan et al., 2019) implementation of a No-U-Turn Sampler (NUTS, Hoffman and Gelman (2011)). We use four chains with 1500 steps (500 burn-in), and additionally average over 10 likelihood realizations per step to aid convergence. A similar analysis was conducted in Pandya et al. (2025) using halotools-IA and IAEmu, allowing us to compare diffHOD-IA against both the ground truth halotools-IA and its emulator IAEmu. HMC on diffHOD-IA took roughly 5 minutes to converge on a single NVIDIA A100 GPU, while HMC with IAEmu converged in approximately one minute on the same GPU. As halotools-IA is not differentiable, a comparable MCMC analysis required up to a full day across 150 CPU cores.
Table 2: Comparison of $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$ posteriors from the Monte Carlo experiments. diffHOD-IA and IAEmu posteriors used HMC, while halotools-IA posteriors used MCMC.
Method $\mu_{\text{cen}}$ $\mu_{\text{sat}}$
halotools-IA $0.793\pm 0.017$ $0.294\pm 0.056$
IAEmu $0.799\pm 0.022$ $0.317\pm 0.049$
diffHOD-IA $0.802\pm 0.016$ $0.318\pm 0.048$
Results. Posteriors for diffHOD-IA, halotools-IA, and IAEmu for this experiment are shown in Figure 6. There is in general excellent overlap between the three posteriors, illustrating that diffHOD-IA offers similar inference capabilities as halotools-IA with much faster convergence with HMC. A slight bias is seen in the case of $\mu_{\text{sat}}$, likely propagated from discrepancies at small scales due to satellite occupation. Nonetheless, diffHOD-IA is within $0.4\sigma$ agreement with halotools-IA. Exact posterior values are given in Table 2.
7 Summary & Discussion
In this work, we have developed diffHOD-IA, a differentiable HOD framework that includes galaxy IA modeling. Our HOD implementation closely follows that of Horowitz et al. (2022), which is extended to the IA modeling of Van Alfen et al. (2024) via inverse Dimroth–Watson CDF sampling. We additionally draw inspiration from Hearin et al. (2022) to extend the differentiability of diffHOD-IA to include IA correlation functions. The diffHOD-IA code is publicly available at .
Speed. The utility of diffHOD-IA over halotools-IA stems from its differentiability, which enables the use of gradient-based algorithms to optimize the input model parameters. For forward modeling, diffHOD-IA runs in comparable time to halotools-IA (seconds per catalog on CPU), with no significant speedup on GPU. For forward modeling, IAEmu is substantially faster, exhibiting an approximately $10000\times$ speedup over halotools-IA and diffHOD-IA. The benefits of differentiability are clearer in inverse modeling. Our HMC analysis converged in approximately 5 minutes on a single GPU, compared to approximately one day across 150 CPU cores for halotools-IA with MCMC.
The utility of diffHOD-IA over IAEmu comes from its differentiability at the catalog level. IAEmu models the correlations $\xi(r)$, $\omega(r)$, and $\eta(r)$, bypassing the galaxy catalog generation step. This limits its predictive abilities to only 2PCFs, whereas diffHOD-IA can be extended to differentiably model any summary statistic. This also requires that any changes to the HOD formulation require retraining IAEmu, while these changes, if necessary, can be made directly within the diffHOD-IA simulation. In addition, IAEmu may not generalize well to different dark matter catalogs (e.g., different cosmologies), whereas this is not a concern with diffHOD-IA.
diffHOD-IA is written in JAX (Bradbury et al., 2018), enabling several processing and vectorization speedups via jax.jit and jax.vmap. As vectorization on GPU requires static array sizes, modeling up to the 2PCF level can be easily vectorized. This does not include generating galaxy catalogs, as the sizes of galaxy catalogs vary across HOD instances and random seeds. For potentially different galaxy-halo connections, or with using weighted-galaxy catalogs, static galaxy catalogs can be generated in parallel.
Results. We benchmarked the accuracy of the sampling procedure and gradients of diffHOD-IA, comparing autodiff computed gradients with analytic and finite-differences, and comparing the differentiable Dimroth Watson samples with samples from the true distribution. diffHOD-IA agreed with halotools-IA across all tests: including galaxy number counts, and the $\xi(r)$, $\omega(r)$, and $\eta(r)$ statistics for the fiducial galaxy catalog.
For a fixed HOD, we utilized gradients in diffHOD-IA to retrieve IA parameters $\bm{\mu}$ from a mock observable galaxy catalog and correlations corresponding to the tng300 simulation. We tested this using gradient-based optimization with a moment matching and correlation function loss. We further leveraged the differentiability of diffHOD-IA in a Hamiltonian Monte Carlo pipeline, showing excellent agreement with halotools-IA and IAEmu.
Future Work. Future extensions of this work could proceed in several directions. The current implementation uses the Zheng et al. (2007a) HOD formulation, but the differentiable framework readily accommodates more sophisticated HOD models that include assembly bias (Hearin and others, 2016) or environment-dependent effects (Yuan et al., 2018). In addition, halotools-IA specifies galaxy misalignments solely according to galaxy orientations, and does not currently include galaxy shapes or ellipticities. diffHOD-IA can readily be extended with shape information in accordance with future versions of halotools-IA. Similarly, while we have focused on the radial alignment model with constant alignment strength, the distance-dependent alignment strength model implemented in diffHOD-IA could be explored for galaxies with radially-varying alignment properties. More generally, this can be extended to different IA parameters per galaxy, as opposed to catalog-wide definitions.
The differentiability also allows inserting diffHOD-IA into larger differentiable inference pipelines. For example, diffHOD-IA can be integrated into simulation pipelines that incorporate a differentiable particle mesh solver such as jaxPM (Li et al., 2022), along with a differentiable halo finder like jFoF (Horowitz et al., 2022). This would enable end-to-end gradient flow from cosmological, HOD, and IA parameters to the galaxy field.
The differentiable correlation function framework could also be extended to 2D-projected statistics, which are more directly comparable to observational weak lensing measurements (Van Alfen et al., 2025a). Of equal interest is the extension to higher order statistics and field-level inference methods which would impose tighter constraints on parameters. A joint inference over both HOD and IA parameters represents a natural application, using the differentiable estimators for $\xi(r)$ and $\omega(r)$. Such joint analyses over the full 7-dimensional parameter space would be more computationally demanding and are reserved for future work. These developments provide a foundation for efficient, gradient-based joint inference of HOD and IA parameters in current and future weak lensing surveys.
Acknowledgments
We thank Nicholas Van Alfen, Alyssa Cordero, Hamza Ahmed, Robin Walters, Andrew Hearin, and Benjamin Horowitz for useful conversations in developing this work. The computations in this paper were run on the FASRC Cannon cluster supported by the FAS Division of Science Research Computing Group at Harvard University. This work is supported by NASA through a Roman Research and Support Participation program (NASA grant 80NSSC24K0088) and the OpenUniverse effort (JPL Contract Task 70-711320); by the DOE (DE-SC0024787); and by the NSF (AST2206563, AST2442796).
Appendix A Computational Efficiency
Figure 7: Runtime in milliseconds for the combined satellite sampling and orientation assignment pipeline as a function of $N_{\rm max}$ and $N_{\rm Newton}$. The red box indicates the fiducial configuration for both parameters used throughout this work. Runtime scales approximately linearly with $N_{\rm max}$, while dependence on $N_{\rm Newton}$ is weaker. Benchmarks use 10,00010{,}000 synthetic host halos and 50,00050{,}000 satellite orientation samples, averaged over 50 runs.
We benchmark the computational performance of diffHOD-IA by comparing the relative costs associated with $N_{\rm max}$, used in the differentiable satellite occupation, and $N_{\rm Newton}$, used for inverse CDF sampling from the Dimroth–-Watson distribution. This analysis focuses on the isolated cost of these computations, not the full diffHOD-IA pipeline. We use a synthetic dataset consisting of 10,000 host halos and 50,000 satellite orientation samples, averaged over 50 runs, executed on a 2021 M1 Max MacBook Pro. The results are shown in Figure 7.
We find that satellite sampling dominates the computational cost, with runtime scaling approximately linearly with $N_{\rm max}$, as illustrated by the color gradient in Figure 7. Although increasing $N_{\rm max}$ improves the validity of the Binomial–Poisson approximation (Equation 15), our fiducial choice of $N_{\rm max}=48$ yields excellent agreement with halotools-IA, as quantified by the galaxy clustering correlation function in Figure 1. As similarly noted by Horowitz et al. (2022), increasing $N_{\rm max}$ beyond this value does not improve the accuracy of relevant summary statistics.
We additionally investigate how varying $N_{\rm Newton}$ impacts the accuracy of the differentiable Dimroth–Watson procedure. Using the fiducial alignment parameters introduced earlier, we compute the percent error in the empirical expectation value of the misalignment angle, $\langle\cos^{2}\theta\rangle$, relative to a high-precision reference computed with $N_{\rm Newton}=50$. We find that the bias decreases with increasing $N_{\rm Newton}$: at $N_{\rm Newton}=2$, the bias is 4.91%, dropping to 0.0365% at $N_{\rm Newton}=4$, and further to $3.4\times 10^{-9}$% at $N_{\rm Newton}=6$. For $N_{\rm Newton}\geq 8$, the bias becomes numerically indistinguishable from zero.
Appendix B Differentiable Galaxy Clustering Statistics
Figure 8: Validation of differentiable correlation function estimators against halotools reference implementations for the same galaxy catalog. Left: Galaxy position-position correlation function $\xi(r)$. The black circles show the (non-differentiable) measurement from halotools.mock_observables.tpcf, while red squares show the differentiable estimator using HOD-derived occupation probability weights. Center: Galaxy position-orientation correlation function $\omega(r)$, comparing halotools.mock_observables.ed_3d (black) with our differentiable estimator (red). Right: Galaxy orientation-orientation correlation function $\eta(r)$, comparing halotools.mock_observables.ee_3d (black) with our differentiable estimator (red). Both $\omega(r)$ and $\eta(r)$ show excellent agreement between implementations, with $\eta(r)$ exhibiting larger statistical fluctuations due to galaxy shape noise. The differentiable estimators enable gradient-based inference: $\xi(r)$ gradients flow through galaxy occupation weights from HOD parameters, while $\omega(r)$ and $\eta(r)$ gradients flow through orientation vectors from IA parameters.
2PCF calculations require galaxy pair counting, which is a discrete operation and not easily made differentiable. To compute 2PCFs differentiably, we follow the methodology of Hearin et al. (2022). This requires treating galaxies as existing with a probability $p$, as opposed to discrete objects with weight 1 (exists) or 0 (does not exist).
For central galaxies, the weight $w_{i}^{\rm cen}$ corresponds to the mean central occupation probability from Equation 2:
$w_{i}^{\rm cen}=\langle N_{\rm cen}(M)\rangle,$ (39)
evaluated at the host halo mass $M$. For satellite galaxies placed in subhalos, the weight is computed as follows:
$w_{i}^{\rm sat}=q_{i}\cdot\langle N_{\rm sat}(M)\rangle,$ (40)
where $q_{i}$ is the softmax weight prioritizing massive subhalos, defined in Equation 16. These weights encode the probability that each galaxy exists in the catalog, enabling gradient flow through the HOD parameters to clustering statistics like $\xi(r)$. Orientation-dependent statistics $\omega(r)$ and $\eta(r)$ receive gradients through the IA parameters.
$\xi(r)$ quantifies the excess probability of finding galaxy pairs at separation $r$ relative to a uniform random distribution:
$\xi(r)=\frac{DD(r)}{RR(r)}-1,$ (41)
where $DD(r)$ is the (weighted) count of galaxy pairs separated by distance $r$, and $RR(r)$ is the expected pair count if galaxies were drawn from a uniform distribution. $\xi(r)$ depends only on the HOD parameters that govern galaxy occupation; the IA parameters $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$ do not affect $\xi(r)$, as they influence only galaxy orientations.
We pre-compute all galaxy neighbor pairs $(i,j)$ with separations $|\mathbf{r}_{ij}|<r_{\rm max}$ using a KD-tree with periodic boundary conditions. The weighted pair count in each bin is computed as:
$DD(r_{k})=\sum_{(i,j)\in\mathcal{B}_{k}}w_{i}\,w_{j},$ (42)
where $w_{i}$ and $w_{j}$ are the galaxy occupation weights and $\mathcal{B}_{k}$ denotes pairs with separations falling in bin $k$. For the expected pair count, we use the analytic expectation:
$RR(r_{k})=\left[\left(\sum_{i}w_{i}\right)^{2}-\sum_{i}w_{i}^{2}\right]\frac{V_{k}}{V_{\rm box}},$ (43)
where $V_{k}=\frac{4\pi}{3}(r_{k+1}^{3}-r_{k}^{3})$ is the shell volume and $V_{\rm box}$ is the simulation volume. The term in brackets represents the effective number of distinct galaxy pairs for the weighted galaxy catalog.
While galaxy positions are fixed in SubhaloPhaseSpace, the weights $w_{i}$ are differentiable functions of the HOD parameters through $\langle N_{\text{cen}}\rangle$ and $\langle N_{\text{sat}}\rangle$. Gradients thus flow through the weight computation:
$\frac{\partial\xi}{\partial\beta}=\sum_{i}\frac{\partial\xi}{\partial w_{i}}\frac{\partial w_{i}}{\partial\beta},$ (44)
where $\beta\in\{\log M_{\rm min},\sigma_{\log M},\log M_{0},\log M_{1},\alpha\}$. Satellites placed via NFWPhaseSpace fallback are assigned weights proportional to $\langle N_{\text{sat}}\rangle$ of their host halo, preserving gradient flow. This enables gradient-based inference of HOD parameters from galaxy clustering measurements.
We benchmark the accuracy of all differentiable correlation functions, including the IA correlations $\omega(r)$ and $\xi(r)$ on the fiducial HOD used in this work. This is shown in Figure 8. There is exact agreement between the diffHOD-IA and halotools-IA estimators for $\omega(r)$, $\eta(r)$, and $\xi(r)$ in the case of unit galaxy weights. In the case of HOD weights, which is necessary for full differentiability of $\xi(r)$ with respect to HOD parameters, the two correlations more noticeably differ, but still exhibit good agreement.
Appendix C Optimization Hyperparameters
We provide details on the optimization hyperparameters used in the gradient-based experiments of Section 6.
C.1 Moment-Matching Optimization
For the moment-matching objective described in Section 6.1, we optimize using the Adam optimizer (Kingma and Ba, 2017) with an initial learning rate of $5\times 10^{-2}$ and an exponential decay schedule (decay rate 0.95 applied every 20 steps). We run the optimization for 100 steps. Gradients are clipped to a global norm of 10.0 to ensure numerical stability. Parameters are initialized by sampling $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$ uniformly from [1,1][-1,1]. We repeat the optimization procedure 50 times with different random initializations.
C.2 Correlation Function Optimization
For the correlation function objective described in Section 6.2, we optimize $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$ using the Adam optimizer with an exponentially decaying learning rate schedule. The initial learning rate is set to $5\times 10^{-3}$, with a decay factor of 0.85 applied every 100 steps. We run the optimization for 2000 steps. Gradients are clipped to a global norm of 1.0 to prevent instabilities from large gradient magnitudes. Parameters are constrained to $|\mu|<1$ via clipping after each update.
The target correlation $\hat{\omega}(r)$ is computed by averaging over 20 orientation realizations of the fiducial tng300 catalog, providing a low-noise reference. We use 20 logarithmically-spaced radial bins spanning $0.1<r<16\;h^{-1}\mathrm{Mpc}$. The optimization catalog is generated with a different random seed than any of the target realizations, ensuring that we fit the statistical correlation structure rather than matching galaxy-by-galaxy. As with the moment-matching case, we perform 50 independent optimization runs from random initializations uniformly sampled from [1,1][-1,1].